Compute both t-tests on excel type a conclusion for each problem

Mr. V. wants to know if grades for his classes have changed since they went online. To test for it, he decided to compute an independent samples t-test and a related samples t-test.

1) For an independent sample, he compared grades for his class on social psychology with grades for his class on abnormal psychology. These were the results:

Social psychology

Abnormal psychology

89

75

91

57

86

64

90

55

100

74

98

63

78

87

84

81

89

53

93

40

2) For a related sample, he compared grades for his introduction to psychology class before and after they went online.

Students

Before online

After online

1

78

93

2

85

86

3

56

88

4

74

89

5

43

96

6

67

91

7

47

87

8

58

99

9

64

79

10

70

83
Answer #1

1.

Suppose, random variables X and Y denote grades in social psychology and abnormal psychology respectively.

Here, two different groups of students are used to collect data from two different classes. Further we do not know population standard deviation (or variance). So, we have to perform two sample t-test.

We have to test for null hypothesis \tiny H_0:\mu_x=\mu_y

against the alternative hypothesis \tiny H_1:\mu_x\neq\mu_y

Corresponding calculations in EXCEL are as follows.

1596081759766_169_01.png

We reject our null hypothesis if \tiny \text{p-value}<\alpha , level of significance.

We generally test for level of significance 0.10, 0.05, 0.01 or something like these.

So, we reject our null hypothesis.

Hence, based on the given data we can conclude that there is significant evidence that grades are different in social psychology and abnormal psychology.

2.

Here, in two different instances (before and after online classes) grades were observed for same set of students. So, we have to perform paired t-test.

Suppose, random variables X and Y denote grades before and after online classes respectively. Also, random variable D(=Y-X) denotes difference in grades.

We have to test for null hypothesis \tiny H_0:\bar d=0

against the alternative hypothesis \tiny H_1:\bar d\neq 0

Corresponding calculations in EXCEL are as follows.

phpTH2b61.png

We reject our null hypothesis if \tiny \text{p-value}<\alpha , level of significance.

We generally test for level of significance 0.10, 0.05, 0.01 or something like these.

So, we reject our null hypothesis.

Hence, based on the given data we can conclude that there is significant evidence that grades are different before and after online classes.